OpenStudy (anonymous):

In the below system, solve for y in the first equation. x - 3y = -6 2x - 7y = 10 A) 1/3x + 2 B) -1/3x + 6 C) -x + 2 D) -1/3 x + 2

OpenStudy (wolf1728):

x -3y = -6 -3y = -6-x y= (-6-x) / -3 y = 2 + (x/3) so answer is A

OpenStudy (anonymous):

awh thank you, and i actually understood how you did it.... something not common for me aha thanks @wolf1728